# Exploring Exotic Decays of the Higgs Boson to Multi-Photons at the LHC via Multimodal Learning Approaches

### v1

@@ 1 Introduction
Another mystery in the SM is the origin of accidental global $U(1)$ symmetries of baryon and lepton numbers which may suggest the presence of new physics beyond the SM. A widely discussed proposal involves elevating these accidental global $U(1)$ symmetries to gauged ones  [13, 14, 15]. Furthermore, the $U(1)_{B}$ gauge symmetry finds application in the $E_{6}$ grand unified theory [16, 17, 18]. Consequently, the leptophobic $Z^{\prime}_{B}$ model emerges as one of the well-motivated $U(1)^{\prime}$ models beyond the SM, particularly those incorporating DM candidates. Examples of UV completions for such models, with or without DM candidates, are detailed in Refs. [19, 20, 21]. The scenario featuring light (baryonic) DM, in particular, stands out as one of the promising facets for DM models involving the leptophobic $Z^{\prime}_{B}$ [22, 23]. In this study, our focus is on a simplified model featuring $U(1)_{B}$ gauge symmetry, assuming that both the leptophobic $Z^{\prime}_{B}$ and the new scalar boson are much lighter than the Higgs boson. Specifically, the new scalar boson primarily decays into a pair of sub-GeV $Z^{\prime}_{B}$ particles, with each $Z^{\prime}_{B}$ predominantly decaying into a neutral pion and a photon, if kinematically allowed.
@@ 2.1 The model
The model with $U(1)_{B}$ gauge symmetry is one of the well-motivated $U(1)^{\prime}$ models beyond the SM [19, 20, 21]. In this simplified model, we assume that all exotic new particles (exotic fermions for anomaly cancellations and extra scalar bosons that provide masses to the exotic fermions and/or make exotic charged/colored particles decay fast enough) are at a relatively high energy scale, except for the light leptophobic $Z^{\prime}_{B}$ and new scalar boson $h_{1}$. The associated $U(1)_{B}$ charges $Q^{i}_{B}$ for fields in our simplified model are listed in Table 1222Note that the Dirac (Majorana) fermion DM candidate $\chi_{D}$ ($\chi_{M}$) with $Q^{i}_{B}(\chi_{D,M})=1$ can be involved in this simplified leptophobic $Z^{\prime}_{B}$ model. Since the DM part is not the focus of this study, we will address this issue in the future..
@@ 4.1 Events generation and detector simulation
To study these signal signatures, we establish a benchmark point (BP) for TP1 and four BPs for TP2, all with fixed values of $\alpha_{B}=10^{-5}$, $M_{Z^{\prime}_{B}}=0.4$ GeV, and $\sin\theta_{h}=1.6\times 10^{-4}$. It’s worth noting that for these BPs, $BR(h_{2}\rightarrow h_{1}h_{1})$ is approximately $8.70\times 10^{-4}$, and $BR(h_{2}\rightarrow Z^{\prime}_{B}Z^{\prime}_{B})$ is roughly $4.35\times 10^{-4}$. UFO model file is generated from FeynRules [72] for the simplified leptophobic $Z^{\prime}_{B}$. Madgraph5 aMC@NLO [73, 74] is used to compute the total cross section and event generation. Generated events are showered by Pythia8 [75] to model the parton showering and hadronization. Delphes3 [76] with modifed ATLAS template is used for the fast detecor simulation. FastJets [77] is used to cluster collimated photons in the final state and apply the basic photon isolation criteria in Delphes3, allowing to distinguish between $\gamma_{\text{rec}}$ and $J_{\gamma}$. The results are summarized in Table 2 and the left panel of Fig. 5, with the average number of reconstructed photons denoted as $\langle N_{\gamma}\rangle$. For TP1, the dominant classification is $\gamma_{\text{rec}}\gamma_{\text{rec}}$ while TP2–BP1 and TP2–BP2 exhibit sizable probabilities for $\gamma_{\text{rec}}J_{\gamma}$. In contrast, TP2–BP3 and TP2–BP4 are primarily classified as $J_{\gamma}J_{\gamma}$.
@@
TP1 -- $75.56\%$ $21.70\%$ $2.75\%$ $1.65$
TP2–BP1 $1.0$ $5.55\%$ $31.18\%$ $63.27\%$ $0.40$
TP2–BP2 $1.2$ $1.65\%$ $18.71\%$ $79.64\%$ $0.21$
TP2–BP3 $1.4$ $4.97\times 10^{-3}$ $9.49\%$ $90.02\%$ $0.10$
TP2–BP4 $1.6$ $2.58\times 10^{-3}$ $4.78\%$ $94.96\%$ $0.05$
Table 2: Branching fractions for $\gamma_{\text{rec}}\gamma_{\text{rec}}$, $\gamma_{\text{rec}}J_{\gamma}$, $J_{\gamma}J_{\gamma}$ and the number of reconstructed photons, $\langle N_{\gamma}\rangle$ for the five BPs with fixed $\alpha_{B}=10^{-5}$, $M_{Z^{\prime}_{B}}=0.4$ GeV and $\sin\theta_{h}=1.6\times 10^{-4}$
Branching fractions for the number of reconstructed photons corresponding to these BPs in the right panel of Fig. 5, for further confirmation of this observation. , for these BPs, is also provided in Table 2. In the following a detailed analysis, particularly regarding the $\gamma_{\text{rec}}J_{\gamma}$ and $J_{\gamma}J_{\gamma}$ signatures, are considered.
Figure 5: The branching fractions for $\gamma_{\text{rec}}\gamma_{\text{rec}}$, $\gamma_{\text{rec}}J_{\gamma}$, and $J_{\gamma}J_{\gamma}$ (left panel), as well as the number of reconstructed photons, $\langle N_{\gamma}\rangle$, (right panel) for the five BPs shown in Table 2.
@@ 4.2 Preselection and event reconstruction
where $\epsilon^{\text{iso}}$ represents the efficiency for collimated photons to satisfy the photon isolation criteria by ATLAS  [26]; more details about this criteria discussed later. The $\epsilon^{\text{ID}}$ follows Eq. (1) in Ref. [40], where "ID" refers to photon identification based on various shower shape variables in different calorimeter layers. According to Fig. 7 in Ref. [40], we estimate $\epsilon^{\text{ID}}_{\text{BSM}}/\epsilon^{\text{ID}}_{\text{SM}}\sim 0.7$ for this study. Although, $\epsilon^{\text{ID}}_{\text{BSM}}$ for TP2 should be smaller than the estimated value, the exact value can still computed using the GEANT4 detector simulation [79], which is beyond the scope of this work.
We first define the combinations as follows: "TP1+TP2-BP1" as Scenario I, "TP1+TP2-BP2" as Scenario II, "TP1+TP2-BP3" as Scenario III, and "TP1+TP2-BP4" as Scenario IV. Next, we present $\epsilon^{\text{iso}}_{\text{TP1}}$, $\epsilon^{\text{iso}}_{\text{TP2}}$, and $(\sigma\times B_{\gamma\gamma})_{\text{BSM}}$ based on five BPs in Table 2 for these four scenarios:
  • Scenario I: $\epsilon^{\text{iso}}_{\text{TP1}}=45.34\%$, $\epsilon^{\text{iso}}_{\text{TP2-BP1}}=3.13\%$, $(\sigma\times B_{\gamma\gamma})_{\text{BSM}}=5.62$ fb;
  • Scenario II: $\epsilon^{\text{iso}}_{\text{TP1}}=45.34\%$, $\epsilon^{\text{iso}}_{\text{TP2-BP2}}=7.39\times 10^{-3}$, $(\sigma\times B_{\gamma\gamma})_{\text{BSM}}=5.10$ fb;
  • Scenario III: $\epsilon^{\text{iso}}_{\text{TP1}}=45.34\%$, $\epsilon^{\text{iso}}_{\text{TP2-BP3}}=1.09\times 10^{-3}$, $(\sigma\times B_{\gamma\gamma})_{\text{BSM}}=4.97$ fb;
  • Scenario IV: $\epsilon^{\text{iso}}_{\text{TP1}}=45.34\%$, $\epsilon^{\text{iso}}_{\text{TP2-BP4}}=6.60\times 10^{-5}$, $(\sigma\times B_{\gamma\gamma})_{\text{BSM}}=4.97$ fb.
@@ 5 Transformer network
Transformer models, introduced by Ref. [43], have revolutionized the field of deep learning and introduced in HEP in Refs. [86, 87, 42, 88, 89] . Unlike traditional sequence-based models, transformers utilize multi-heads self-attention mechanism to process input data in parallel, making them highly efficient for capturing long-range dependencies. This inherent capability makes transformers promising candidates for analyzing complex and interconnected systems, such as particle clouds. Adapting transformer models to analyze particle clouds involves encoding the spatial and temporal dynamics inherent in event collisions. Representing each particle as an input token, transformers can effectively capture interactions and correlations between particles. Moreover, self-attention mechanism allows the model to weigh the significance of each particle’s contribution to the overall cloud behavior. In this manner, transformer model can capture the important features shared by each individual particle to the whole cloud allowing for better classification performance between different signal and background events.
@@ 6.1 Projected limits on the model parameters
Regarding the systematic uncertainties arising from QCD di-jet measurements, which may affect the reported classification performance, we consider a $20\%$ systematic uncertainty to estimate this effect. The formula for signal significance is modified accordingly:
@@ 6.2 Decoding networks output
Figure 12 visualizes the attention scores, $A_{i,j}$, generated by the additive attention layer for $10^{5}$ signal events (left panel) and background events (right panel). Intriguingly, the network predominantly focuses on the first five highest momentum particles inside the leading and subleading jet to identify signal events, whereas for background events, the network distributes high attention across a wider range of particle contents. We interpret this behavior as the network learning the distinct substructures of photon and QCD jets.

### v2

@@ 1 Introduction
Another mystery in the SM is the origin of accidental global $U(1)$ symmetries of baryon and lepton numbers which may suggest the presence of new physics beyond the SM. A widely discussed proposal involves elevating these accidental global $U(1)$ symmetries to gauged ones  [13, 14, 15, 16]. Furthermore, the $U(1)_{B}$ gauge symmetry finds application in the $E_{6}$ grand unified theory [17, 18, 19]. Consequently, the leptophobic $Z^{\prime}_{B}$ model emerges as one of the well-motivated $U(1)^{\prime}$ models beyond the SM, particularly those incorporating DM candidates. Examples of UV completions for such models, with or without DM candidates, are detailed in Refs. [20, 14, 21, 15, 16, 22]. The scenario featuring light (baryonic) DM, in particular, stands out as one of the promising facets for DM models involving the leptophobic $Z^{\prime}_{B}$ [23, 24]. In this study, our focus is on a simplified model featuring $U(1)_{B}$ gauge symmetry, assuming that both the leptophobic $Z^{\prime}_{B}$ and the new scalar boson are much lighter than the Higgs boson. Specifically, the new scalar boson primarily decays into a pair of sub-GeV $Z^{\prime}_{B}$ particles, with each $Z^{\prime}_{B}$ predominantly decaying into a neutral pion and a photon, if kinematically allowed.
@@ 2.1 The model
The model with $U(1)_{B}$ gauge symmetry is one of the well-motivated $U(1)^{\prime}$ models beyond the SM [20, 14, 21, 15, 16, 22]. In this simplified model, we assume that all exotic new particles (exotic fermions for anomaly cancellations and extra scalar bosons that provide masses to the exotic fermions and/or make exotic charged/colored particles decay fast enough) are at a relatively high energy scale, except for the light leptophobic $Z^{\prime}_{B}$ and new scalar boson $h_{1}$. The associated $U(1)_{B}$ charges $Q^{i}_{B}$ for fields in our simplified model are listed in Table 1222Note that the Dirac (Majorana) fermion DM candidate $\chi_{D}$ ($\chi_{M}$) with $Q^{i}_{B}(\chi_{D,M})=1$ can be involved in this simplified leptophobic $Z^{\prime}_{B}$ model. Since the DM part is not the focus of this study, we will address this issue in the future..
@@ 4.1 Events generation and detector simulation
To study these signal signatures, we establish a benchmark point (BP) for TP1 and four BPs for TP2, all with fixed values of $\alpha_{B}=10^{-5}$, $M_{Z^{\prime}_{B}}=0.4$ GeV, and $\sin\theta_{h}=1.6\times 10^{-4}$. It’s worth noting that for these BPs, $BR(h_{2}\rightarrow h_{1}h_{1})$ is approximately $8.70\times 10^{-4}$, and $BR(h_{2}\rightarrow Z^{\prime}_{B}Z^{\prime}_{B})$ is roughly $4.35\times 10^{-4}$. UFO model file is generated from FeynRules [73] for the simplified leptophobic $Z^{\prime}_{B}$. Madgraph5 aMC@NLO [74, 75] is used to compute the total cross section and event generation. Generated events are showered by Pythia8 [76] to model the parton showering and hadronization. Delphes3 [77] with modified ATLAS template is used for the fast detecor simulation. Moreover, we have implemented the photon conversion module in Delphes3, parameterized according to the conversion map of the CMS tracker as shown in Fig. 1 of Ref. [78]. Note that the magnetic field in the detector is also simulated in the photon conversions module, which has been improved from the simulation in Ref. [31].FastJets [79] is used to cluster collimated photons in the final state. The basic photon isolation criteria in Delphes3 and the photon isolation criteria used by ATLAS [27] [^5] Details about the ATLAS photon isolation criteria are provided in Sec. 4.2. are applied, allowing for the distinction between $\gamma_{\text{rec}}$ and $J_{\gamma}$.
The results are summarized in Table 2, with the average number of reconstructed photons denoted as $\langle N_{\gamma}\rangle$. The results without the photon conversion simulation are also shown in brackets for comparison. We find that the effect of photon conversions appears limited due to the following two main reasons: (1) The multiple collimated photons from the Higgs boson via gluon fusion are not energetic enough, resulting in $e^{+}e^{-}$ pairs that are too soft; and (2) soft objects generated from Pythia8 simulations, such as multiparton interactions (MPI), have already significantly affected photon isolation. Additionally, we note that the photon transverse shower is not simulated in this analysis. In Refs. [27, 31], the photon transverse shower is considered a powerful variable for distinguishing between photons and QCD jets. However, since the current version of Delphes3 for fast detector simulation does not include the simulation of the photon transverse shower, we did not include it in our analysis [^6] Additionally, simulating the Moliere radius [31] could further improve the accurate identification of the number of photons within the photon jet, but this is beyond the scope of this work.. For TP1, the fraction of each classification is approximately one-third while TP2–BP1 and TP2–BP2 exhibit sizable probabilities for $\gamma_{\text{rec}}J_{\gamma}$. In contrast, TP2–BP3 and TP2–BP4 are primarily classified as $J_{\gamma}J_{\gamma}$. In the following, a detailed analysis is conducted, particularly regarding the $\gamma_{\text{rec}}J_{\gamma}$ and $J_{\gamma}J_{\gamma}$ signatures.
@@
TP1 -- $27.51\%$ ($32.84\%$) $36.36\%$ ($35.11\%$) $36.13\%$ ($32.05\%$) $0.91$ ($1.01$)
TP2–BP1 $1.0$ $4.44\%$ ($5.53\%$) $30.59\%$ ($34.58\%$) $64.97\%$ ($59.89\%$) $0.39$ ($0.46$)
TP2–BP2 $1.2$ $2.69\%$ ($2.94\%$) $23.36\%$ ($27.11\%$) $73.95\%$ ($69.95\%$) $0.29$ ($0.33$)
TP2–BP3 $1.4$ $1.12\%$ ($1.36\%$) $17.52\%$ ($19.50\%$) $81.36\%$ ($79.14\%$) $0.20$ ($0.22$)
TP2–BP4 $1.6$ $0.59\%$ ($0.77\%$) $13.06\%$ ($14.05\%$) $86.35\%$ ($85.18\%$) $0.14$ ($0.16$)
Table 2: Branching fractions for $\gamma_{\text{rec}}\gamma_{\text{rec}}$, $\gamma_{\text{rec}}J_{\gamma}$, $J_{\gamma}J_{\gamma}$ and the number of reconstructed photons, $\langle N_{\gamma}\rangle$ for the five BPs with fixed $\alpha_{B}=10^{-5}$, $M_{Z^{\prime}_{B}}=0.4$ GeV and $\sin\theta_{h}=1.6\times 10^{-4}$. The numbers in brackets here are without the photon conversion simulation.
@@ 4.2 Preselection and event reconstruction
where $\sigma=48.52$ pb is the Higgs boson production via gluon fuison at $\sqrt{s}=13$ TeV with N3LO QCD corrections and NLO electroweak corrections [80]. $\epsilon^{\text{iso}}$ represents the efficiency for collimated photons to satisfy the photon isolation criteria by ATLAS [27]; more details about this criteria discussed later. The $\epsilon^{\text{ID}}$ follows Eq. (1) in Ref. [41], where "ID" refers to photon identification based on various shower shape variables in different calorimeter layers. According to Fig. 7 in Ref. [41], we estimate $\epsilon^{\text{ID}}_{\text{BSM}}/\epsilon^{\text{ID}}_{\text{SM}}\sim 0.7$ for this study. Although, $\epsilon^{\text{ID}}_{\text{BSM}}$ for TP2 should be smaller than the estimated value, the exact value can still computed using the GEANT4 detector simulation [81], which is beyond the scope of this work.
We first define the combinations as follows: “TP1+TP2-BP1” as Scenario I, “TP1+TP2-BP2” as Scenario II, “TP1+TP2-BP3” as Scenario III, and “TP1+TP2-BP4” as Scenario IV. Next, we present $\epsilon^{\text{iso}}_{\text{TP1}}$, $\epsilon^{\text{iso}}_{\text{TP2}}$, and $(\sigma\times B_{\gamma\gamma})_{\text{BSM}}$ based on five BPs in Table 2 for these four scenarios:
  • Scenario I: $\epsilon^{\text{iso}}_{\text{TP1}}=27.51\%$, $\epsilon^{\text{iso}}_{\text{TP2-BP1}}=4.44\%$, $(\sigma\times B_{\gamma\gamma})_{\text{BSM}}=3.76$ fb;
  • Scenario II: $\epsilon^{\text{iso}}_{\text{TP1}}=27.51\%$, $\epsilon^{\text{iso}}_{\text{TP2-BP2}}=2.69\%$, $(\sigma\times B_{\gamma\gamma})_{\text{BSM}}=3.40$ fb;
  • Scenario III: $\epsilon^{\text{iso}}_{\text{TP1}}=27.51\%$, $\epsilon^{\text{iso}}_{\text{TP2-BP3}}=1.12\%$, $(\sigma\times B_{\gamma\gamma})_{\text{BSM}}=3.08$ fb;
  • Scenario IV: $\epsilon^{\text{iso}}_{\text{TP1}}=27.51\%$, $\epsilon^{\text{iso}}_{\text{TP2-BP4}}=0.59\%$, $(\sigma\times B_{\gamma\gamma})_{\text{BSM}}=2.97$ fb.
@@ 5 Transformer network
Transformer models, introduced by Ref. [44], have revolutionized the field of deep learning and introduced in HEP in Refs. [88, 89, 43, 90, 91] . Unlike traditional sequence-based models, transformers utilize multi-heads self-attention mechanism to process input data in parallel, making them highly efficient for capturing long-range dependencies. This inherent capability makes transformers promising candidates for analyzing complex and interconnected systems, such as particle clouds. Adapting transformer models to analyze particle clouds involves encoding the spatial and temporal dynamics inherent in event collisions. Representing each particle as an input token, transformers can effectively capture interactions and correlations between particles. Moreover, self-attention mechanism allows the model to weigh the significance of each particle’s contribution to the overall cloud behavior. In this manner, transformer model can capture the important features shared by each individual particle to the whole cloud allowing for better classification performance between different signal and background events. We point out to previous photon jets analysis using MLP, CNN and GNN [92, 93, 94], which utilize image or graph-based input datasets. An improved representation for these analyses is an unordered, permutation-invariant set of particles, known as a particle cloud. This representation retains all the benefits of particle-based models, especially the flexibility to incorporate arbitrary features for each particle. Similar to the point cloud representation used in computer vision, particle clouds offer a robust framework for modeling jets. Unlike graph or image representations that require particles to be sorted and thus implicitly depend on particle order, particle clouds impose no such order. Since particles within a jet lack intrinsic order, any manually imposed arrangement may be suboptimal and negatively impact performance. Additionally, particle clouds do not require preprocessing steps, which are essential for image and graph-based datasets and can cause networks to heavily rely on these transformations. Traditional networks such as MLPs, CNNs, or basic GNNs are not permutation invariant and therefore cannot effectively analyze particle cloud datasets. In contrast, the Transformer network is permutation invariant and well-suited for particle cloud analysis. Additionally, ATLAS experiment utilize the photon transverse energy deposited in the ECAL to distinguish photons and misidentified hadronic jets [95]. This variable has a discriminating validity as the energy clusters deposited by photons in the EM calorimeter tend to be narrower in the transverse direction than those deposited by jets.
@@ 6.1 Projected limits on the model parameters
Regarding the systematic uncertainties arising from QCD di-jet measurements, which may affect the reported classification performance, we consider a $20\%$ systematic uncertainty to estimate this effect. The formula for signal significance is modified accordingly [96, 97]:
@@ 6.2 Decoding networks output
Figure 11 visualizes the attention scores, $A_{i,j}$, generated by the additive attention layer for $10^{5}$ signal events (left panel) and background events (right panel). Intriguingly, the network predominantly focuses on the first five highest momentum particles inside the leading and subleading jet to identify signal events, whereas for background events, the network distributes high attention across a wider range of particle contents. We interpret this behavior as the network learning the distinct substructures of photon and QCD jets. Therefore, it makes the network robust against random fluctuations in the photon jet substructure caused by potential pileup effects.